In a Math test, the mean score is 48 and the standard deviation is 7. Assuming normality, what is probability that a score picked at random will lie above score 55?

1
Expert's answer
2022-04-11T17:19:38-0400

"P(X>55)=P(Z>\frac{55-48}{7})=P(Z>1)=1-P(Z<1)=0.1587."